How do you solve y=5x+11
y=-4x-7?
I've tried substitution and many other ways to solve the problem but it's complicated.

Found 2 solutions by Alan3354, jsmallt9:
Answer by Alan3354(69443)

How do you solve y=5x+11
y=-4x-7?
I've tried substitution and many other ways to solve the problem but it's complicated.
--------------
Both RH sides = y, so they equal each other.
5x+11 = -4x-7
9x = -18
x = -2
-------
y = 1

Answer by jsmallt9(3759)

y = 5x + 11
y = -4x - 7
There are many methods to solve this. When deciding which method to use, it is helpful to know how to pick a method which will be easy. For the substitution method, look for a variable that is already "solved for" or can easily be "solved for". ("Solved for" means the variable is by itself on one side of the equation.)

Since the "y" is already "solved for" in the first equation (and the second, for that matter), the Substitution method will probably be easiest method.

With this method, after you have solved for one of the variables in one of the equations, you substitute for that variable in the other equation. Since y is already "solved for" in the first equation we can substitute for y in the other equation. Replacing the y in the second equation with the 5x + 11 from the first equation we get:
5x + 11 = -4x - 7

We now have an equation of one variable. This is relatively easy to solve. Adding 4x to each side we get:
9x + 11 = -7
Subtracting 11 from each side we get:
9x = -18
Dividing by 9 we get:
x = -2

It is tempting to stop here. But we need both the x and the y of the solution. The next step is to use the value we got for x and substitute it into one of the original equations and solve for y. Substituting the -2 for x in the first equation we get:
y = 5(-2) + 11
Simplifying...
y = -10 + 11
y = 1

So the solution to this system is the point (-2, 1)